OpenStudy (cypress7):

Cybil read 3/10 h before breakfast, 9/20 h during the day, and 1 3/5 h after dinner? How many total hours did Cybil read? Express the answer in simplest form.

OpenStudy (cypress7):

@ganeshie8 @SolomonZelman @dan815

OpenStudy (mathstudent55):

You need to add the numbers: \(\dfrac{3}{10} + \dfrac{9}{20} + 1 \dfrac{3}{5} \)

OpenStudy (cypress7):

Okay thank you

OpenStudy (mathstudent55):

You're welcome.